# Short-term plasticity recalls forgotten memories through a trampoline mechanism

### v1

@@ C.1 Final self-consistent DMFT with plastic kernel
Appendix D Numerical solution of the DMFT equations
In this Appendix we describe the numerical procedure used to obtain self-consistent solutions of the DMFT equations derived in Appendix C. The method follows the standard approach used in dynamical mean-field analysis of disordered systems [eissfeller1992new, eissfeller1994mean, roy2019numerical]: we replace the continuous-time formulation by a discretized representation on a finite time grid, express all temporal kernels as matrices acting on this grid, and iterate the resulting equations until convergence. This numerical scheme provides an explicit realization of the effective stochastic process governing a representative neuron in a network of infinitely large size, together with the associated response and correlation functions.
D.1 Time discretization
@@
The effective single-site dynamics derived in Appendix C is recapitulated here again for completeness:
@@
which represents the cumulative influence of past activity through the effective feedback kernel $M$. On the discretized time grid, this convolution is approximated by a left Riemann sum[hughes2020calculus]. To simplify the notation, we introduce the matrix $M_{ij}\equiv M(t_{i},t_{j})$, which remains strictly lower-triangular because the continuous kernel $M(t,\tau)$ vanishes for $t<\tau$ by causality. The integral $I(t)$ becomes then
@@
D.2 Single-site response and $R_{\phi}$
@@
The resulting matrix $\mathbf{R}_{\phi}$ is strictly lower-triangular (up to numerical noise), consistent with causality, and in practice we enforce this structure by projecting out any spurious entries above the diagonal. The matrix $\mathbf{R}_{\phi}$ is then used as the continuous $O(1)$ kernel in the construction of the Hopfield response $R_{m}$ and the corresponding noise covariance, as discussed in Appendix D.3.
D.3 Discretization of $R_{m}$ and the Hopfield kernel
@@
D.4 Noise covariance
In Appendix C we derived the continuous-time expression for the noise covariance, recapitulated here as
@@ D.4 Noise covariance
This expression is the discrete analogue of the continuous convolution structure: the noise at two times $t_{i}$ and $t_{k}$ is correlated whenever the uncondensed modes can jointly propagate fluctuations from earlier times $t_{i}$ and $t_{j}$. We rewrite this expression using the normalized operators normalization introduced in Sec. D.3, $(\mathbf{R}_{m}^{\mathrm{op}})_{i,j}=R_{m}(t_{i},t_{j})\,\Delta t$, and obtain
@@ D.5 Main algorithm
    Forward integration. For each realization $r$ and each time index $i=1,\dots,T-1$, we integrate the discretized dynamics using the Euler scheme discussed in Sec. D.1, namely
@@
Appendix E Finite size simulations
The finite size simulations presented in Fig. 3 are obtained by solving numerically Eqs. 1 and 5, for a system of $N=20000$ neurons, using Euler algorithm with a time-step $\Delta t=0.25$. The dynamical overlap is computed from the average $m^{*}(t)=\frac{1}{N}\sum_{i=1}^{N}\langle\xi^{*}_{i}\phi(x_{i}(t))\rangle$, where the average $\langle\cdot\rangle$ is obtained by averaging over $2048$ replicates with different realizations of the target overlap. The target pattern is conventionally chosen to be the first one generated when constructing the matrix $J_{ij}$, $\xi^{*}_{i}\equiv\xi^{1}_{i}$. The network is initialized using the same procedure described in Eqs. 129 and 130. We show in Fig. 6 the temporal evolution of the overlap $m^{*}(t)$ at large times for a finite-size system.
Figure 6: Long-time dynamics of the neural network without short-term synaptic plasticity We show the temporal evolution of the overlap $m^{*}(t)$against a target pattern for a network of $N=20000$ neurons with $k=0$.
@@ Appendix E Finite size simulations
Appendix F Jacobian stability analysis
@@
It is possible to show that $\lambda$ is a root of (147) if and only if $\lambda$ is an eigenvalue of $\mathbf{M}_{\text{comp}}$.

### v2

@@ D.1 Final self-consistent DMFT with plastic kernel
Appendix E Numerical solution of the DMFT equations
In this Appendix we describe the numerical procedure used to obtain self-consistent solutions of the DMFT equations derived in Appendix D. The method follows the standard approach used in dynamical mean-field analysis of disordered systems [12, 13, 47]: we replace the continuous-time formulation by a discretized representation on a finite time grid, express all temporal kernels as matrices acting on this grid, and iterate the resulting equations until convergence. This numerical scheme provides an explicit realization of the effective stochastic process governing a representative neuron in a network of infinitely large size, together with the associated response and correlation functions.
E.1 Time discretization
@@
The effective single-site dynamics derived in Appendix D is recapitulated here again for completeness:
@@
which represents the cumulative influence of past activity through the effective feedback kernel $M$. On the discretized time grid, this convolution is approximated by a left Riemann sum[26]. To simplify the notation, we introduce the matrix $M_{ij}\equiv M(t_{i},t_{j})$, which remains strictly lower-triangular because the continuous kernel $M(t,\tau)$ vanishes for $t<\tau$ by causality. The integral $I(t)$ becomes then
@@
E.2 Single-site response and $R_{\phi}$
@@
The resulting matrix $\mathbf{R}_{\phi}$ is strictly lower-triangular (up to numerical noise), consistent with causality, and in practice we enforce this structure by projecting out any spurious entries above the diagonal. The matrix $\mathbf{R}_{\phi}$ is then used as the continuous $O(1)$ kernel in the construction of the Hopfield response $R_{m}$ and the corresponding noise covariance, as discussed in Appendix E.3.
E.3 Discretization of $R_{m}$ and the Hopfield kernel
@@
E.4 Noise covariance
In Appendix D we derived the continuous-time expression for the noise covariance, recapitulated here as
@@ E.4 Noise covariance
This expression is the discrete analogue of the continuous convolution structure: the noise at two times $t_{i}$ and $t_{k}$ is correlated whenever the uncondensed modes can jointly propagate fluctuations from earlier times $t_{i}$ and $t_{j}$. We rewrite this expression using the normalized operators normalization introduced in Sec. E.3, $(\mathbf{R}_{m}^{\mathrm{op}})_{i,j}=R_{m}(t_{i},t_{j})\,\Delta t$, and obtain
@@ E.5 Main algorithm
    Forward integration. For each realization $r$ and each time index $i=1,\dots,T-1$, we integrate the discretized dynamics using the Euler scheme discussed in Sec. E.1, namely
@@
Appendix F Finite size simulations
The finite size simulations presented in Fig. 3 are obtained by solving numerically Eqs. 1 and 5, for a system of $N=20000$ neurons, using Euler algorithm with a time-step $\Delta t=0.25$. The dynamical overlap is computed from the average $m^{*}(t)=\frac{1}{N}\sum_{i=1}^{N}\langle\xi^{*}_{i}\phi(x_{i}(t))\rangle$, where the average $\langle\cdot\rangle$ is obtained by averaging over $2048$ replicates with different realizations of the target overlap. The target pattern is conventionally chosen to be the first one generated when constructing the matrix $J_{ij}$, $\xi^{*}_{i}\equiv\xi^{1}_{i}$. The network is initialized using the same procedure described in Eqs. 129 and 130. We show in Fig. 6 the temporal evolution of the overlap $m^{*}(t)$ at large times for a finite-size system, both with and without short-term synaptic plasticity.
Figure 6: Finite size simulations of the long-time dynamics of the neural network. We show the temporal evolution of the overlap $m^{*}(t)$against a target pattern for a network of $N=20000$ neurons for different values of the short-term synaptic plasticity strength $k$.
@@ Appendix F Finite size simulations
Appendix G Jacobian stability analysis
@@
It is possible to show that $\lambda$ is a root of Eq. (147) if and only if $\lambda$ is an eigenvalue of $\mathbf{M}_{\text{comp}}$.